A car travels 300 Km in the same time that a freight train travels 200 Km. The speed of the car is 20 Km/h more than the speed of the train. Find the speed of the car and the speed of the train.
Answer by mananth(16949)

Freight train x 200 miles
Car x 20 300 miles

t=d/r
200/x=300/(x+20)
Cross multiply
200(x+20)=300*x
200x+4000=300x
300x-200x=4000
100x=4000
/100
x=40Freight train mph
60 mph Car